OpenStudy (anonymous):

In the 3rd quadrant , if sin A = -3/5 and sec B = -13/12 then tan (A-B) = ?

OpenStudy (anonymous):

sinA=-3/5 secB=-13/12(I think you know that secB is the same as 1/cosB) I'll draw the triangles for you..

OpenStudy (anonymous):